OpenStudy (anonymous):

A mobile game uses two servers to handle matchmaking: one on the West Coast and one on the East Coast. The East server has 50 free users and 20 premium users, while the West server has 40 free users and 40 premium users. While the process for selecting the next user is random, there is a preference given for premium users: on each pick there is a 60% than a premium user is selected and a 40% chance a free user is selected. What is the probability that the next player will be a free user from the East server?

OpenStudy (anonymous):

OMG those eyes! ima have a nightmare today D:

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

this is tough question

OpenStudy (anonymous):

^^; any idea how to do it?

OpenStudy (anonymous):

I'm pretty lost.

OpenStudy (nirmalnema):

from east coast. 40% of 50 = 20 free users.. and 60% of 20 is 12 random user.. from west coast.. 40% of 40 is 16 free user and 60% of 40 is 24 random users total users = 20+24+16+12=72 free users from east coast.=20 so its probablity is 72/20 36/10=3.6 ...

OpenStudy (nirmalnema):

may be but im not fully sure

OpenStudy (anonymous):

ur ans is nice ur right its not sure

OpenStudy (nirmalnema):

yup..

OpenStudy (anonymous):

Thank you for the help anyways! : )

OpenStudy (nirmalnema):

welcome..!!

OpenStudy (nirmalnema):

do u understand what i have done

OpenStudy (anonymous):

I get what you were trying to do : )

OpenStudy (nirmalnema):

ok..